OpenStudy (anonymous):

The period of the graph of the function f(x)=2tan2x is___? a. pi b. pi/2 c. pi/4 d. 2pi and why?

OpenStudy (anonymous):

It's b because the period of tan(x) is Pi and the factor 2 makes it half.

OpenStudy (anonymous):

b. pi/2

OpenStudy (anonymous):

How do we know that the period of tan(x) is pi? Is it already a given?

OpenStudy (anonymous):

Yes. You can also derive it from \[ \tan(x)=\frac{\sin(x)}{\cos(x)} \]